OpenStudy (anonymous):

How can I rationalize (r^1/2 - r^-1/2)^2 ?

OpenStudy (mertsj):

\[(\sqrt{r}-\frac{1}{\sqrt{r}})^2=r-2+\frac{1}{r}\]

OpenStudy (blockcolder):

\[\large \left (\sqrt{r}-\frac{1}{\sqrt{r}}\right )^2=\left ( \frac{r-1}{\sqrt{r}}\right )^2=\frac{r^2-2r+1}{r}=r-2+\frac{1}{r}\]

OpenStudy (mertsj):

Use FOIL

OpenStudy (anonymous):

Thank you :)

OpenStudy (blockcolder):

No problem. :D